OpenStudy (princessaurora):

*Trigonometric Identities* Given tan(x)=1/3 and cos x < 0, find the value of cos(x)? *Will fan and give medal*

OpenStudy (princessaurora):

help por favor

OpenStudy (pawanyadav):

You can use Sec^2 x =1+tan^2 x Then put sec^2 x=1/(cos^2 x)

OpenStudy (princessaurora):

Is that the answer?

OpenStudy (pawanyadav):

Its Hint

OpenStudy (princessaurora):

I don't know how to do it

OpenStudy (pawanyadav):

Is answer {-3/√10}

OpenStudy (green_1):

is that the answer @Pawanyadav

OpenStudy (princessaurora):

GRACIAS

OpenStudy (pawanyadav):

Use above identity Put tan x=1/3 Sec ^2 x= 1+(1/3)^2 I/(cos^2 x)=10/9 So cos^2 x=9/10 Then cos x<0 So cos x=- √{9/10}=-3/√10

OpenStudy (princessaurora):

What is sec(x)?

OpenStudy (pawanyadav):

1/cos x=sec x=-(√10)/3

OpenStudy (princessaurora):

Thank you so much

OpenStudy (pawanyadav):

Welcome